How fast is the size of the population changing after 1 minute? 


1
Expert's answer
2021-10-29T03:02:03-0400
"P(t)=3t^2+3e^t+10, t\geq0"

"P'(t)=(3t^2+3e^t+10)'=6t+3e^t"

Given "t=1\ min=60\ sec."


"P'(60)=6(60)+3e^{60}=360+e^{60}\approx1.142\times10^{26}"

The popukation increases at rate of "360+e^{60}\approx1.142\times10^{26}" bacteria per second after 1 minute.